# Using iterated local alignment to aggregate trajectory data into a traffic flow map

### v2

@@ 3.3 Line blending to align similar linestrings with local reference
11:         $\mathcal{F}:={\tt ST\_OVERLINE\_LINEBLEND}(\mathcal{F},k,\varepsilon,\varepsilon_{S})$
12:         $\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
13:$\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
14:while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
15:     $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
16:     $\mathcal{F}:={\tt ST\_OVERLINE\_PRUNE}(\mathcal{F})$
4 Results
In this section we compute a minimal flow map for the Hannover trajectories. From the complete set of 1183 trajectories, we keep 1177 trajectories with length greater than $100$ m. We input the 1177 trajectories into ST_ROUTE with $M$ as the map matching and $R$ the route finding APIs from the Valhalla routing engine, and ${\bm{n}}_{W}=3,13,23,33,43,63,83$ waypoints. We employ the dockerised version 3.4.0 of the Valhalla routing engine (GIS OPS, 2023). Of these input trajectories, 1147 yield a sufficiently high quality match, where the Hausdorff distance $d_{\operatorname{Haus}}(M^{*}(G),G)<100$ m, and the ratio $\operatorname{len}(M^{*}(G))/\operatorname{len}(G)<1.1$. We continue the analysis with these 1147 map matched routes.
For all iterations, we set $j_{\max}=20$. For the line blending, we begin with 1 iteration of ST_OVERLINE (Steps 1–4) with simplify tolerance $\varepsilon_{D}=1~{}\mathrm{m}$. We follow with ST_OVERLINE (Steps 5–14) with blend and snap tolerances $\varepsilon=\varepsilon_{S}=4~{}\mathrm{m}$, #edges ${\bm{k}}=1,2$, and node split type $S=$ ‘subdivision’. These simplify, blend and snap tolerances were chosen heuristically as a trade-off between being sufficiently large to account for the noisiness of the trajectories and the map matching/route finding APIs, whilst not being too large to obscure separate road segments within region with a dense road network. We set ${\bm{k}}=1,2$ since the search for connected $k$-edges with $k>2$ is too computationally intensive for our setup due to the large number of road segments (13 495). We set $S=$ ‘subdivision’ node splitting, as it provides more stable line blending priority at this early stage. We follow with 1 iteration of ST_OVERLINE (Steps 5–14), with the same tuning parameter choices, except with $S=$ ‘unary’. ‘Unary’ node splitting is usually applied after an iteration of ‘subdivision’ node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority. We end with 2 iterations of ST_OVERLINE (Steps 5–14), with $\varepsilon=\varepsilon_{S}=5~{}\mathrm{m},{\bm{k}}=1,2,3,4,S=$ ‘unary’, since the larger values for $\varepsilon,\varepsilon_{S}$ capture any line blending missed at 4 m, and the connected 3-, 4-edge searches become computationally feasible with the lower number of road segments. These 5 iterations of ST_OVERLINE converge to a minimal flow map with 1 413 road segments, i.e. a 89.5% reduction of the 13 495 segments in the initial flow map.
Figure 9 illustrates the results of these iterations. In Figure 9(a) are the trajectories (green circles) and their map matched routes resulting from ST_ROUTE (blue lines). Whilst the map matched routes no longer completely obscure the road network, they remain misaligned to each other and to the road network. So it is not possible to accurately estimate the traffic flow map directly from them. In Figure 9(b) is a flow map resulting from ST_OVERLINE (Steps 1–4) with $S=$ ‘subdivision’. This is similar to the flow map that would be obtained by following Morgan and Lovelace (2021) without rasterisation. Since each linestring is labelled by its flow value, the crowding of the labels indicates that this is unlikely to be a minimal flow map. In Figure 9(c), we complete the line blending iteration ST_OVERLINE (Steps 5–14) with $S=$ ‘subdivision’. The crowding of the flow value labels is reduced so we are progressing to a minimal flow map. In Figure 9(d), we carry out a further 3 line blending iterations ST_OVERLINE (Steps 5–14) with $S=$ ‘unary’ to arrive at a minimal flow map.
Refer to caption(a) Refer to caption(b)
Refer to caption(c) Refer to caption(d)
Figure 9: Iterations to compute locally aligned flow map. (a) Empirical trajectories (green circles), map matched routes from ST_ROUTE (blue lines). (b) Initial flow map from ST_OVERLINE (1–4), $S=$ ‘subdivision’. (c) Intermediate flow map from 1 iteration of ST_OVERLINE (5–14), $S=$ ‘subdivision’. (d) Minimal flow map from 3 further iterations of ST_OVERLINE (5–14), $S=$ ‘unary’. Label is traffic flow. Colour (purple to orange) of road segments is proportional to traffic flow.
4.1 Validation
A visual inspection of the minimal flow map in Figure 9(d) reveals good alignment to the OSM road network in general, though there remain some data artefacts elsewhere. For example, in Figure 10(a), the green trajectories give no indication of the blue wonky route output by ST_ROUTE, and in Figure 10(c), the trajectories do not involve a loop, though ST_ROUTE contains a loop. Since our local alignment does not use an external road network to correct these wonky or extraneous loops in the map matched routes, they are propagated to the flow maps in Figure 10(a, d). Overall, these type of data artefacts are small and infrequent, and are associated with road segments with low flow.
Refer to caption(a) Refer to caption(b)
Refer to caption(a) Refer to caption(b)
Figure 10: Data artefacts in locally aligned road segment flows. (a–b) Wonky road segments. (c–d) Extraneous road segments. (a, c) Trajectories (green circles), map matched routes (blue lines). (b, d) Colour (purple to orange) of road segments is proportional to traffic flow.
For a more quantitative validation of the accuracy of our proposed flow map, we require a gold standard reference flow map. The experimental design of the Hannover trajectories is to serve as a reference data set for learning turning rules at road junctions, rather than as a reference flow map (Zourlidou et al., 2022). So we have to compute a proxy reference flow map. For this, we first compute line transects $\mathcal{T}_{i}=\{{\bm{t}}_{i,1},\dots,{\bm{t}}_{i,n_{\mathcal{T}_{i}}}\}={\tt ST\_TRANSECT}({\bm{f}}_{i},\varepsilon_{T},\delta_{T})$ for a road linestring ${\bm{f}}_{i}\in\mathcal{F}$, where a line transect is an orthogonal line segment (of length $2\varepsilon_{T}$ m) to ${\bm{f}}_{i}$. If $\operatorname{len}({\bm{f}}_{i})>\delta_{T}$ m, then these transects are placed at every $\delta_{T}$ m of ${\bm{f}}_{i}$, or if $\operatorname{len}({\bm{f}}_{i})\leq\delta_{T}$ m then the single transect is subtended from the centroid of ${\bm{f}}_{i}$. The empirical proxy flows are the number of intersection points between the map matched routes from $\mathcal{M}$ which intersect the line transects $f_{i,j}^{\mathrm{emp}}=\#\{{\tt ST\_INTERSECTION}(\mathcal{M},{\bm{t}}_{i,j}):{\bm{t}}_{i,j}\in\mathcal{T}_{i}\},j=1,\dots,n_{\mathcal{T}_{i}}$, and the mean discrepancy over all line transects of ${\bm{f}}_{i}$ is $\bar{f}_{i}^{\mathrm{emp}}=\mathrm{mean}\{f_{i,j}^{\mathrm{emp}},j=1,\dots,n_{\mathcal{T}_{i}}\}$. Our error measure of ${\bm{f}}_{i}$ is the absolute difference $\operatorname{Err}_{i,j}=|f_{i}-\bar{f}_{i}^{\mathrm{emp}}|$ for all line transects ${\bm{t}}_{j},j=1,\dots,n_{\mathcal{T}_{i}}$ which intersect ${\bm{f}}_{i}$. We also compute the relative error $\operatorname{RErr}_{i,j}=\operatorname{Err}_{i,j}/f_{i}$, which is well-defined since $f_{i}\geq 1$ for all ${\bm{f}}_{i}\in\mathcal{F}$.
With the tuning parameters tolerance $\varepsilon_{T}=5$ m and $\delta_{T}=50$ m, then we compute 7 380 line transects. The colour (purple to orange) of the line transects is proportional to the absolute error. In Figure 11(a), the vast majority of the error values are low (purple), with only a few high error values (orange), in/near the solid black rectangle. The zoom of the black rectangle is given in Figure 11(b), along with the road segment flows in blue. The road segment ${\bm{f}}_{11}$ (thick blue segment) with the estimated flow is $f_{11}=238$, and the empirical proxy flow $\bar{f}_{11}^{\mathrm{emp}}=254$, has a high absolute error, i.e. $\operatorname{Err}_{11}=16,\operatorname{RErr}_{11}=0.07$. The road segment ${\bm{f}}_{331}$ with $f_{331}=25,\bar{f}_{331}^{\mathrm{emp}}=12$, has a higher relative error, i.e. $\operatorname{Err}_{331}=13,\operatorname{RErr}_{331}=0.52$. These errors are due to that these road segments, at an earlier iteration of the flow map, comprised shorter sub-segments with higher flows due to the presence of the more than $100$ trajectories passing near the intersection of ${\bm{f}}_{11}$ and ${\bm{f}}_{331}$. Since the nodes of these sub-segments are pseudo nodes, they were removed, and the concatenated sub-segments assigned the single weighted mean flow = 25.
Refer to caption(a) Refer to caption(b)
Figure 11: Validation of line transects of locally aligned flow map. (a) Neighbourhood level. (b) Zoom of black rectangle. Line transects computed with tolerance $\varepsilon_{T}=5$ m, at every $\delta_{T}=50$ m. Label is absolute error (Err). Colour (purple to orange) of line transects is proportional to absolute error. Width of road segments (blue) is proportional to traffic flow.
To supplement the visual examination of these flow estimation errors, Figure 12 shows the bivariate (Err, RErr) histogram plot of the errors. The vast majority (6518 or 88.3%) in the orange hexagonal bin have zero absolute and relative error, and only 39 (0.53%) have Err $>4$, and 364 (4.9%) have RErr $>0.1$, and only 16 (0.21%) have Err $>4$ and RErr $>0.1$ which comprise most of the purple bins.
Figure 12: Histogram plot of absolute and relative errors of line transects of locally aligned flow map. Label in hexagonal bin is number of line transects. Vertical dotted line is Err = 4, horizontal dotted line is RErr = 0.1.
This demonstrates that our flow map is mostly accurate as it usually controls the estimation error. By accuracy, we mean by how close the estimated flows are to the empirical proxy flows. However accuracy is insufficient on its own: if we take the extreme situation of a flow map with a single road segment with zero error, then this has the highest accuracy, but we have no estimated flows outside of this single segment. So we require that the spatial coverage of the flow map be also high. We cannot resolve this question of spatial coverage unambiguously since we do not have a gold standard flow map, though we can verify that all $n=1147$ map matched routes intersect at least one line transect from the estimated flow map. Whilst this calculation does not exclude that there can be some regions of some map matched routes are without nearby line transects, we can further verify with a visual inspection that these regions tend to be small in area and/or comprise a low number of routes. So we claim that our minimal flow map has high levels of accuracy and spatial coverage.
To conclude the validation of our flow maps, the flow map at the city level is illustrated in Figure 13(a). The orange segments with high flows are apparent, whereas these high flows were not apparent from the scatter plot of the trajectories in Figure 1(a). The desire line map (or spider diagram) is in Figure 13(b) and represent the traffic flows between the origin/destination hub nodes (black solid circles) These hub nodes are the centroids from a hierarchical clustering, similar to that in ST_SNAPNODE, of the boundary points of trajectories with the cutoff at 5 000 m. The straight lines indicate the trajectories whose origin/destination are associated with different hub nodes, and the circles indicate the trajectories whose origin/destination are associated with the same hub node. The desire line map is in effect a low resolution map of straight-line flows between the hub nodes only, whereas the high resolution map shows the traffic flows on all road segments.
Refer to caption(a) Refer to caption(b)
Figure 13: High and low resolution flow maps. (a) Locally aligned flow map. (b) Desire line map. Colour (purple to orange) and width of road segments is proportional to traffic flow.
4.2 Software
These analysis algorithms have been developed in R since the complexity of Algorithms 18 require a mix of advanced geospatial and statistical methods. As a compiled language, R can have slower execution times. There are two main computational bottlenecks. The first consists of the map matching/route finding in Algorithm 1. The Valhalla routing engine APIs are available as a web-based service (e.g. https://valhalla1.openstreetmap.de) or as a local dockerised image (e.g. https://github.com/gis-ops/docker-valhalla). We use the latter as it allows for faster computation since these local API requests are not sent to a remote web-based server, and can be parallelised on a stand-alone machine. We conduct a small study of the execution times based on 10 replicates of ST_ROUTE on 10 randomly selected trajectories on an Intel i5 Quad core 3.10 GHz machine running Ubuntu 22.04 and R 4.4.0. Executing ST_ROUTE with a local Valhalla API is around 7.1 times faster than the web-based API, and a parallelisation (with 3 workers) is around 1.8 times faster than a serial computation. This is less than 3 because the dockerised image is not optimised for simultaneous API calls. Combining these together, a parallelised local API achieves around a 12.6 fold speed improvement in comparison to a serial web-based API.
The second bottleneck is the line blending in Algorithm 7. Based on the execution times based on 10 replicates of ST_OVERLINE on a subset of 433 trajectories, 3-worker parallelisation is around 1.8 times faster than a serial computation. This is less than 3 because only the repeated application of line blending (Algorithm 4) is parallelised, whilst the line blending priority (Algorithm 6) remains a serial computation. These speed factors are intended to be illustrative, since execution times, involving remote web servers APIs and parallelisation, are difficult to predict on different internet connections and machines. We tentatively claim that a local Valhalla API reduces the execution time by an order of magnitude, whereas parallelisation reduces it almost linearly by the number of workers.
Our long term goal is to integrate fast implementations of these local alignment algorithms into an industry standard GIS, so GIS practitioners will be able to efficiently create flow maps from empirical trajectories in a familiar environment. Since this is under development, in the mean time, we provide a geopackage and QGIS project with the input trajectories, map matched routes, iterated flow maps and desire lines, as listed in Table 1. The interested reader is able to interactively explore in QGIS the added value of our proposed high resolution minimal flow map flowmap4, in comparison to the input trajectories traj, the map matched routes route, the desire line flow map flowmap_desire, and the flow map computed according to a leading alternative without rasterization Morgan and Lovelace (2021) which is similar to flowmap0.
Layer Description $n$
traj Empirical trajectories 1 147
route Map matched routes ST_ROUTE 1 147
flowmap0 Flow map ST_OVERLINE(1–4), $\varepsilon_{D}=1~{},S=$ ‘subdivision’ 13 495
flowmap1 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,\varepsilon=\varepsilon_{S}=4,S=$ ‘subdivision’ 2 437
flowmap2 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,\varepsilon=\varepsilon_{S}=4,S=$ ‘unary’ 1 953
flowmap3 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,3,4,\varepsilon=\varepsilon_{S}=5,S=$ ‘unary’ 1 867
flowmap4 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,3,4,\varepsilon=\varepsilon_{S}=5,S=$ ‘unary’ 1 413
flowmap_desire Desire lines ST_DESIRELINE, $\varepsilon=5000$ 41
Table 1: Geospatial layers in geopackage. The first column is layer name, the second is description, and the third is number of geospatial features $n$.

### v5

@@ 2.4 Line blending to align similar linestrings with local reference
11:     $\mathcal{F}:={\tt ST\_OVERLINE\_PRUNE}(\mathcal{F})$
2.5 Workflow
Our proposed workflow from an input of trajectory data to a locally aligned flow map is illustrated in Figure 6. We begin with the trajectories $\mathcal{G}$. If the measurement errors in these trajectories are small enough to be input directly into ST_OVERLINE, then we set $\mathcal{R}=\mathcal{G}$. Otherwise to obtain $\mathcal{R}$, we carry out map matching using ST_ROUTE for each number of waypoints ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value. For the flow map aggregation, we begin with an iteration of ST_OVERLINE_INIITAL with simplify tolerance $\varepsilon_{D}=1$ m to create an initial flow map. In Figure 6, the dotted rectangle surrounding ST_LINEBLEND_PLANR and ST_SNAPNODE indicates that these algorithms are inputs to ST_OVERLINE_INIITAL. For the subsequent steps, we set the maximum number of iterations $j_{\max}=20$, and the snap tolerance $\varepsilon_{S}=\varepsilon$. We call ST_OVERLINE with #edges ${\bm{k}}=1,2$, and node split type $S=$ “subdivision”. At this early stage, the search for connected $k$-edges with $k>2$ can be computationally intensive, and “subdivision” node splitting provides more stable line blending priority. The dotted rectangle indicates the algorithms which are inputs to ST_OVERLINE. We follow with a call to ST_OVERLINE with “unary” node splitting. The “unary” node splitting is usually applied after “subdivision” node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority. After these two steps, the flow map has begun to stabilise, and 3-, 4-edge searches become computationally feasible with the lower number of flow segments. We iterate ST_OVERLINE with ${\bm{k}}=1,2,3,4,S=$ “unary”, until the flow map $\mathcal{F}$ converges. In this workflow, the only unspecified tuning parameter is the blend tolerance $\varepsilon$, whose properties we investigate in the next section.
Figure 6: Workflow for locally aligned flow map. Input are trajectories $\mathcal{G}$, output is locally aligned flow map $\mathcal{F}$. Blend tolerance $\varepsilon$ is tuning parameter.
3 Results and discussion
3.1 Synthetic trajectories
Despite the ubiquity of GNSS-enabled devices, there is a lack of gold standard empirical trajectories and flow maps, due to privacy concerns and other reasons. Our solution is to rely on noise-free synthetic trajectories which are exactly aligned to a theoretical road network (Anastasiou et al., 2022a). These authors generate synthetic trajectories from an origin-destination matrix, and make available a set of 1.5 million synthetic trajectories for the metropolitan area of Los Angeles, USA (Anastasiou et al., 2022b). For illustrative purposes, we focus on a sample of $n=100$ trajectories, in EPSG 6339 (NAD83 (2011)/UTM zone 10N) coordinates, in Culver City (located about 16 km from downtown Los Angeles) which pass near the intersection of Main Street and Venice Boulevard, as indicated by the solid yellow line in Figure 7. We add a random uniform noise of ±$\pm 5$ m to the noise-free trajectories to obtain the noisy trajectories, displayed as the green lines in Figures 7(a–b). This noise mimics the typical measurement error of 5 m observed in empirical trajectories from GNSS-enabled smartphones (Van Diggelen and Enge, 2015). The reference noise-free flow map is in Figure 7(c), and the rasterised flow map in Figure 7(d). Morgan and Lovelace (2021) do not recommend a value for tuning the rasterisation grid resolution, so we select $\varepsilon_{R}=2$ m subjectively so that the raster grid cells do not exceed the road widths.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Synthetic Culver City trajectories $(n=100)$. (a) Noisy trajectories (green lines). (b) Zoom of noisy trajectories near intersection of Main St and Venice Bd (solid yellow line). (c) Reference noise-free flow map. (d) Rasterised flow map with $\varepsilon_{R}=2$ m. Label in flow maps is traffic flow.
For our locally aligned flow maps, the free tuning parameter is the blend tolerance parameter $\varepsilon$. So we investigate the effect of $\varepsilon$ on the accuracy of flow aggregation. For the blend tolerance $\varepsilon=1$ m, the flow map in Figure 8(a) is unable to consistently blend the map matched routes correctly since 1 m is typically smaller than the remaining discrepancies after map matching in ST_ROUTE. For the blend tolerance $\varepsilon=5$ m, the central road segments labelled with flow 51 and 45 in Figure 8(b), which are close to 49 and 45 in the reference flow map. For the blend tolerance $\varepsilon=10$ m, the central road segments are also labelled with flow 50 and 45 in Figure 8(c), similar to with blend tolerance $\varepsilon=5$ m in Figure 8(b). Near the intersection of Venice Bd and Main St (solid yellow line), the flow segment (with flow 17) is poorly aligned to the road network, unlike the more accurate alignment in Figure 8(b).
Refer to caption Refer to caption
Figure 8: Effect of blend tolerance $\varepsilon$ on synthetic Culver City flow maps. (a) Flow map with $\varepsilon=1$ m. (b) Flow map with $\varepsilon=5$ m. (c) Flow map with $\varepsilon=10$ m. Label is traffic flow.
To complement the above heuristic comparisons, we define a quantitative discrepancy between a flow map $\mathcal{F}$ and the reference flow map $\mathcal{F}_{\rm ref}$. First, we compute the transect points for each linestring ${\bm{f}}$ from $\mathcal{F}$: if $\operatorname{len}({\bm{f}})>\varepsilon_{T}$ m, then a point is placed at every $\varepsilon_{T}$ m of ${\bm{f}}$; or if $\operatorname{len}({\bm{f}})\leq\varepsilon_{T}$ m, then a single point is placed at the centroid of ${\bm{f}}$. So $\mathcal{T}=\{{\bm{t}}_{1},\dots,{\bm{t}}_{n_{\mathcal{T}}}\}$ is a set of $n_{\mathcal{T}}$ transect points. For each ${\bm{t}}_{i},i=1,\dots,n_{\mathcal{T}}$, we find the closest linestring in $\mathcal{F}_{\rm ref}$, that is, ${\bm{f}}_{{\rm ref},i}^{*}=\operatorname{argmin}_{{\bm{f}}_{\rm ref}\in\mathcal{F}_{\rm ref}}\mathrm{dist}({\bm{t}}_{i},{\bm{f}}_{\rm ref})$. The flow errors are $\operatorname{Err}_{i}=|f_{i}-f_{{\rm ref},i}^{*}|$ for the transect points on ${\bm{t}}_{i}\in\mathcal{T}\subseteq\mathcal{F}$. We reverse the order of the comparison to obtain the flow errors $\operatorname{Err}_{j}=|f_{{\rm ref},j}-f_{j}^{*}|$ for the transect points on ${\bm{t}}_{j}\in\mathcal{T}_{\rm ref}\subseteq\mathcal{F}_{\rm ref}$, and concatenate these two sets of flow errors.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Flow errors for synthetic Culver City flow maps. (a) Flow map with $\varepsilon=1$ m. (b) Flow map with $\varepsilon=5$ m. (c) Flow map with $\varepsilon=10$ m. (d) Rasterised flow map with $\varepsilon_{R}=2$ m. Flow segments are coloured lines, transect points are grey circles, flow errors are labels.
In Figure 9(a), we display the transect points with $\varepsilon_{T}=50$ m, as the grey circles, superposed over the flow map $\mathcal{F}$. Figures 9(a–c) are the flow errors for the locally aligned flow maps with blend tolerance $\varepsilon=1,5,10$ m, and Figure 9(d) is for the rasterised flow map. The $\varepsilon=1$ m flow map and the rasterised flow map lead to many high error values. Whereas for the other two locally aligned flow maps, the flow errors are similarly low.
We repeat the flow map error calculations for $N=10$ re-samples of $n=100$ Culver City trajectories. The mean and SD of the flow errors for the locally aligned flow maps are $3.061\pm 0.565$ ($\varepsilon=1$ m), $0.446\pm 0.047$ ($\varepsilon=5$ m), $0.477\pm 0.048$ ($\varepsilon=10$ m), and for the rasterised flow map $1.514\pm 0.201$ ($\varepsilon_{R}=2$ m). Setting $\varepsilon=1$ m leads to under-blending since $\varepsilon$ is smaller than the misalignments in the map matched routes, and setting $\varepsilon=10$ m leads to over-blending since $\varepsilon$ is larger than the distance between separate road segments. The $\varepsilon_{R}=2$ m rasterised flow map is similarly under-blended like the $\varepsilon=1$ m flow map. So we tentatively recommend a value of about $\varepsilon=5$ m for the blend tolerance parameter since it leads to optimal line blending.
3.2 Empirical trajectories
We return to the empirical Hannover trajectories. From the 1183 trajectories, we keep 1177 trajectories with length greater than $100$ m, and follow the workflow in Figure 6. We map match these 1177 trajectories with ST_ROUTE, and we continue with 1152 of them which have a sufficiently high quality match, where the Hausdorff distance $\operatorname{dist}_{\operatorname{Haus}}(R(G),G)<100$ m, and the ratio $\operatorname{len}(R(G))/\operatorname{len}(G)<1.1$, for a trajectory $G$ and a map matched route $R(G)$. That is, we discard 25 trajectories or a 2.02% failure rate for the Valhalla map matching, which is consistent with previous studies (Saki and Hagen, 2022).
After some trial and error, we set the blend tolerance of $\varepsilon=4$ m, which is 1 m less than the 5 m tentatively recommended from the synthetic trajectories computations above. The workflow leads to a locally aligned flow map displayed in Figure 10(a). The corresponding rasterised flow map is in Figure 10(b). At the city level, both of these flow maps visualise well the traffic flows. Though if we zoom into Klusreide (solid yellow circle), then we observe that in Figure 10(c) the locally aligned map displays the aggregated flows on linear road segments in comparison to the pixelated road segments of the rasterised flow map in Figure 10(d).
Whilst we do not have a reference flow map for these empirical trajectories, like we had for the synthetic Los Angeles trajectories, we are still able to compute a proxy flows based on empirical counts of the map matched routes with orthogonal line transects from the flow map road segments. Along Klusriede, the locally aligned flow map is composed to two road segments (in pink) with 0 and 2 flow errors) corresponding to the proxy flows 312 and 315. For the rasterised flow map, some grid cells in the pink region have flow error 2. Though rasterisation some times leads to double counting and thus to much larger flow errors, e.g. in the orange region near the T-junction, the estimated flow is 643 whereas the proxy flow is 315, so the flow error is 328. The locally aligned map does not suffer from these rasterisation effects, so its flow aggregation is more accurate.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 10: Hannover flow maps. (a) Locally aligned flow map at city level. (b) Rasterised flow map at city level. Solid yellow circle is Klusriede. (c) Locally aligned flow map at road segment level. (d) Rasterised flow map at road segment level. Colour scale is traffic flow, label is traffic flow error.
Our future goal is the integration of the local alignment algorithms into an industry standard web GIS. In the mean time, we make available interactive web maps at https://mvstat.net/traj which allow the user to freely explore our proposed multi-scale traffic flow maps from the individual road to the city-wide level.
3.3 Computation
We report the execution times for the flow map computations for the synthetic and empirical trajectories. Since it is difficult to reliably reproduce execution times on different systems, we provide our execution times as a guide only. These are obtained on a stand-alone set-up running Ubuntu 24.04 with 16Gb RAM and 4 × Intel i5 CPU @ 3.10 GHz. We deploy a local dockerised Valhalla routing engine API (https://github.com/gis-ops/docker-valhalla) (GIS OPS, 2023) rather than relying on a web-based service. Since the local map matching API requests can be efficiently parallelised on a stand-alone machine, we are confident that ST_ROUTE is suitably optimised.
For $n=100$ synthetic Culver City trajectories, ST_ROUTE executes in 188 s. For $n=1000$ synthetic trajectories, it executes in 2233 s, which is 10 times longer for a 10-fold increase in $n$. So it is straightforward to infer a precise execution time of ST_ROUTE for large sample sizes from a small sample size. The execution time for ST_OVERLINE increases as $n$ increases, but not in the same uniform manner as ST_ROUTE as the former also depends on the value of $\varepsilon$. For $n=100$, the execution times for ST_OVERLINE are 1777 s ($\varepsilon=1$ m), 1049 s ($\varepsilon=5$ m), 880 s ($\varepsilon=10$ m); and for $n=1000$, they are 15 457 s ($\varepsilon=1$ m), 6002 s ($\varepsilon=5$ m), 3897 s ($\varepsilon=10$ m). So it is difficult to infer precisely the execution times for ST_OVERLINE for larger sample sizes and smaller blend tolerances from smaller $n$ and larger $\varepsilon$. This is because the line blending (ST_LINEBLEND) is parallelised, but the line blending priority (ST_LINEBLEND_PRIORITY) remains a serial computation.
With an optimal value of the blend tolerance $\varepsilon=5$ m, the total execution time of the workflow for the $n=100$ Culver City trajectories is 1238 s or $\sim$21 mins, and for $n=1000$ is 8236 s or $\sim$137 mins. The total execution time for the $n=1183$ Hannover trajectories with $\varepsilon=4$ m is 5545 s or $\sim$92 mins. Thus locally aligned flow maps are currently suitable for offline analyses of small and medium sized data sets. Due to time and computing constraints, (a) data sets with a larger number of trajectories (like the full set of 1.5 million synthetic Los Angeles trajectories) and (b) online computations were out of scope. For these to be computationally feasible, we require additional implementation improvements in the local alignment algorithms, which we leave for future development.